The length of a rectangle is 4 less than 3 times the width. If the
perimeter is 56 inches, what is the area of the rectangle?

Answer by waynest(281)

P = Lw
length (L) = 3w - 4
width (w)
Perimeter = 56
2w + 2(3w -4) = 56
2w + 6w - 8 = 56
8w - 8 + 8 = 56 + 8
8w = 64
8w/8 = 64/8
w = 8
length = 3(8) - 4
= 24 - 4
= 20
check:
2(8) + 2(3(8) - 4) = 56
16 + 2(24 - 4) = 56
16 + 2(20) = 56
16 + 40 = 56
56 = 56
Area = Lw
= 20 * 8
= 160 in^2